OpenStudy (lovelycharm):

Printing a newsletter costs $1.50 per copy plus $450 in printer's fees. The copies are sold for $3 each. how many copie of the newsletter must be sold to break even?

OpenStudy (anonymous):

Expenses = f(x) = 450 + 1.50x, where x is the number of newsletters to be printed. Income = g(x) = 3x. To break even, Income - Expenses must equal zero.

OpenStudy (lovelycharm):

@Aeronautix so what the answer is

OpenStudy (anonymous):

g(x) - f(x) = [3x] - [450+1.50x] = 0, then solve for x. That is the number of newsletters that must be sold to break even.

OpenStudy (lovelycharm):

@Aeronautix i got 300 copie